OpenStudy (anonymous):

The manufacturing cost of a cell phone was $136. Sellers first marked up the cost by 25 percent. Sales were poor, so sellers decreased the markup to 17 percent. Calculate the old and new selling prices. selling price = original cost + markup value

OpenStudy (pradipgr817):

old= 136+25% new =old-17%of old

OpenStudy (pradipgr817):

old= 136+25% of 136

OpenStudy (anonymous):

Thannks ")

OpenStudy (anonymous):

The Old Is 170 I Cant Find The New ,